OpenStudy (anonymous):

how would you find the solutions for.. x^4 - 16 = 0

OpenStudy (anonymous):

Add 16 to both sides. Then raise both sides to the 1/4 th power

geerky42 (geerky42):

Don't forgot \(\pm\), though.

OpenStudy (radar):

factor for sure. (x + 2)(x-2)(x+4)=0

geerky42 (geerky42):

Actually it's (x-2) (x+2) (x²+4).

OpenStudy (mathstudent55):

Think of x^4 - 16 = 0 as (x^2)^2 - 4^2 = 0 and factor using the difference of two squares. Then factor the difference of two squares again.

OpenStudy (radar):

I don't believe (x+4) will work x=-4 -4^4 -16=0 ??????

geerky42 (geerky42):

\[(x-2)(x+2)(x^2+4)\]

OpenStudy (radar):

Please disregard my solution as I erred.

OpenStudy (radar):

@geerky42, you are correct it is x^2+4 giving you the complex roots.

geerky42 (geerky42):

So \(\Large x^4 - 16 = (x-2)(x+2)(x-2i)(x+2i) = 0\) So there are four solutions. (two real and two complex) x - 2 = 0 x + 2 = 0 x - 2i = 0 x + 2i = 0 \(x = 2, ~-2, ~2i, ~\text{or}~-2i\)

geerky42 (geerky42):

Is this clear? @McKailaMarie2014

OpenStudy (anonymous):

yes. thank you!

geerky42 (geerky42):

Glad we helped.